Circular No. 5831

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence,

write: “CCD photometry of the optical counterpart of the recurrent x-ray transient V1333 Aql (Aquila X-1), obtained at Haute-Provence with the 1.2-m telescope on July 15, reveals that the source has turned on again and is now at a brightness level (V = 17.7) typical of previous recent episodes. When it was last observed on May 3, the object was still in quiescence. This new event arrives apparently at least two months earlier than expected from predictions based on the 1989-1992 episodes (see IAUC 5665).”

NOVA OPHIUCHI 1993 AND NOVA AQUILAE 1993 R. W. Argyle and L. V. Morrison, Royal Greenwich Observatory, report

accurate positions for N Oph 1993 and N Aql 1993 obtained with the Carlsberg Automatic Meridian Circle on La Palma, operated by J. L. Muinos, D. W. Evans, P. Jensen, and B. Jensen. In both cases, the observed positions are given for equinox J2000.0, FK5 reference frame, and epoch 1993.45: N Oph 1993, α = 17^{h}25^{m}06^{s}.410 ± 0^{s}.006, δ = −23^{o}11′11′′.10 ± 0′′.08 (corresponding 1950.0 position: α = 17^{h}22^{m}04^{s}.396, δ = −23^{o}08′32′′.55); N Aql 1993, α = 19^{h}13^{m}06^{s}.804 ± 0^{s}.003, δ = +1^{o}34′23′′.33 ± 0′′.05 (corresponding 1950.0 position: α = 19^{h}10^{m}34^{s}.712, δ = +1^{o}29′14′′.29).

J. Elias, Cerro Tololo Interamerican Observatory (CTIO), communicates that infrared spectrograms of Nova Aql 1993, taken on July 14.1 UT by D. DePoy, R. Blum, and K. Sellgren (Ohio State University), H. Tirado (CTIO), and himself with the OSU Infrared Imager Spectrometer on the CTIO 4-m telescope show a strong, red continuum in the H and K bands, probably indicative of dust formation. Emission lines of H I, He I 1083.0nm, and O I 1128.7-nm are also present.

SUPERNOVA 1993J IN NGC 3031 Further photoelectric photometry by H. Mikuž, B. Dintinjana, and T.

Zwitter, Ljubljana, Slovenia (cf. IAUC 5796): July 7.88 UT, B = 14.13 ± 0.03, V = 13.15 ± 0.02, R = 12.80 ± 0.03, I = 12.44 ± 0.04; 13.86, B = 13.99 ± 0.02, V = 13.51 ± 0.01, R = 13.01 ± 0.02, I = 12.51 ± 0.01.

1993 July 16 Daniel W. E. Green